a) Suppose that Netflix switches to offering a separate service for TV shows and a separate service for movies. Netflix charges its ??? TV customers price ??? and its ?? movie customers ??. Demand for the products are inter-related. Specifically:

??? = 14 − 2??? − ??

And: ?? = 19 − 3?? − 2???

And the cost function is: ? = 10 + 4(?? + ???)

Identify the optimal quantities (?? and ???) for Netflix.

b) Now let’s contrast the separate pricing with the bundled price for both products. Suppose that there are 2 different types of users Movie-buff and TV-aholic. The number of each type and their willingness to pay are shown:

TV only Movie only Number 300,000 700,000 Type Movie-buff TV-aholic Bundle 12 16 10

Assume that there are no costs.

i. What are the revenues from a single optimal bundle price? (Assume you don’t offer the products separately.)

ii. What are the revenues from selling the products separately if you optimally choose both prices? (Assume you don’t offer the bundle.)
Answer #1

A)

TTVE PTV QTU - C - (14-201- QM) QI - 10 - 4(QM+BT) = 14-QM) Q7 - 2072 -10 – 4 COM+QT) so datu 3 (14-QM) - 4QF - 4.=0. dQT 14-

B) (i) bundling

So if bundle price = 12, then both type will buy

So total revenue = 1000,000*12 = 12000,000

(As all people = 1000,000, will buy the bundle)

If price = 16, then movie buff type will not buy the good,

So TR = 700,000*16 = 11200,000

Thus higher TR, when P of bundle = 12, so optimal bundle price = 12

& TR = 12000,000

(Ii) separate pricing

For TV , if P = 4, both type buy

TR = 1000,000*4 = 4000,000

If P = 10, only TV aholic will buy

So TR = 700,000*10 = 7000,000

Thus higher TR, when P = 10

For Movie, if P = 6, both type will buy

TR = 6*1000,000 = 6000,000

If P = 8, only movie buff type will buy

TR = 8*300,000 = 2400,000

So highly TR if p = 6

So TR = 7000,000 + 6000,000

= 13000,000